OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 5. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (yanasidlinskiy):

Use the log laws. \[\log(ab)=\log(a)+\log(b)\] and \[\log _{a} a=1\] so look at \[\log _{16}(32)=\log _{16}(2 \times 16)=\log _{16}(2)+1\] and \[2=16^{1/4}\] so \[\log _{16}(16\frac{ 1 }{ 4 })=1/4\] So the problem is now \[\frac{ 1 }{ 4 }+1=x +5 \] or, \[\frac{ 5 }{ 4 }=x +5\] Now Solve for x!

OpenStudy (anonymous):

would it be x=-15/4?

OpenStudy (yanasidlinskiy):

Yep!!:)

OpenStudy (anonymous):

awesome! thank you!! :D

OpenStudy (yanasidlinskiy):

No problem!!:) Anytime!!!:)